OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, 1). Which equation defines circle P?

OpenStudy (chaise):

sqrt(8)=(x-2)^2+(y-1)^2 I hope this is right!

OpenStudy (amistre64):

almost

OpenStudy (amistre64):

r^2 = x^2 + y^2 is the basic set up; then modified for central positions

OpenStudy (chaise):

sqrt(8)^2=(x-2)^2+(y-1)^2 Is that better :s?

OpenStudy (amistre64):

nope, your hanging onto that sqrt, its bad baggage to begin with

OpenStudy (amistre64):

r=8; therefore r^2 = 8^2

OpenStudy (chaise):

Tell me how you would go about it then. Sorry for any confusion.

OpenStudy (anonymous):

A cone with a radius of 16 feet and a slant height of 27 feet ?What is the surface area of this cone? Use 22/7 for pi.

OpenStudy (anonymous):

Bridget has a kitchen canister, in the shape of a cylinder, which holds flour. If the canister is 22 centimeters wide and 24 centimeters tall, what is its approximate surface area in terms of pi?

OpenStudy (anonymous):

If a tennis ball has a diameter of approximately 6.7 centimeters, what is its surface area?

OpenStudy (amistre64):

you really should post new questions to a new post; otherwise it just gets buried in the pile